A baker is making cakes.
 She starts with 48 eggs.
 She makes 10 cakes.
 She has 18 eggs left over.
Which equation can be used to find the number of eggs, x, the baker uses in
each cake?
A.
10x +18 = 48
B.
18x +10 = 48
C.
10(x +18) = 48
D.
18(x +10) = 48

Answer by dkppathak(439)

A baker is making cakes.
 She starts with 48 eggs.
 She makes 10 cakes.
 She has 18 eggs left over.
Which equation can be used to find the number of eggs, x, the baker uses in
each cake?
A.
10x +18 = 48
B.
18x +10 = 48
C.
10(x +18) = 48
D.
18(x +10) = 48
option A is right option
10X+18=48
10x=48-18=30
x=30/10=3
he need 3 eggs for each cake . He can make 6 more cake